# Reference-Free Sampling-Based Model Predictive Control

### v1

@@ Abstract
We present a sampling-based model predictive control (MPC) framework that enables emergent locomotion without relying on handcrafted gait patterns or predefined contact sequences. Our method discovers diverse motion patterns, ranging from trotting to galloping, robust standing policies, jumping, and handstand balancing, purely through the optimization of high-level objectives. Building on model predictive path integral (MPPI), we propose a dual-space spline parameterization that operates on position and velocity control points. Our approach enables contact-making and contact-breaking strategies that adapt automatically to task requirements, requiring only a limited number of sampled trajectories. This sample efficiency allows us to achieve real-time control on standard CPU hardware, eliminating the need for GPU acceleration typically required by other state-of-the-art MPPI methods. We validate our approach on the Go2 quadrupedal robot, demonstrating various emergent gaits and basic jumping capabilities. In simulation, we further showcase more complex behaviors, such as backflips, dynamic handstand balancing and locomotion on a Humanoid, all without requiring reference tracking or offline pre-training.
@@ I Introduction
(a) Overview of the framework showing the dual-spline parametrization, noise schedule and reference-free costs.
@@
Figure 1: Overview of the reference-free sampling-based MPC framework (top): our approach enables emergent jumping motion experimentally achieved on the Go2 robot without any guiding reference (bottom).
@@
Online robot control presents a trade-off between solution quality and computational efficiency. While learning-based methods, such as reinforcement learning (RL), can generate awe-inspiring movements on complex robots, they require extensive offline training. They may also fail to generalize or adapt well to new environments or tasks that were not seen during training [kober2013rlsurvey]. Despite ongoing efforts to improve sample efficiency, RL remains orders of magnitude too slow for online adaptation, and simplified models used to accelerate training can hinder sim-to-real transfer [deisenroth2011pilco, zhao2020sim2realreview]. Moreover, RL methods often rely heavily on engineered rewards, such as phase clocks, air-time penalties, or foot-clearance objectives, to enforce structured contact behavior [hwangbo2019legged, Aractingi_2023, ha2025learning].
@@
On the other side, trajectory optimization (TO) methods using gradient-based techniques can find high-quality solutions but are computationally intensive and require accurate gradient information. This may be unavailable or unreliable in contact-rich scenarios, leading most TO methods to use predefined contact sequences [budhiraja2018differential, jalletPROXDDPProximalConstrained2025]. Contact-implicit TO removes the need for predefined contact schedules [posa2014contactImplicit] and has enabled advanced demonstrations of complex whole-body behaviors [neunertWholeBodyNonlinearModel2018, kimContactimplicitModelPredictive2025]. However, many formulations rely on simplified contact models or approximations to remain tractable, which can introduce a mismatch with real dynamics and produce only approximate solutions in practice. These methods also necessitate hand-crafted cost functions to obtain good contact sequences [kimContactimplicitModelPredictive2025].
@@
Sampling-based methods offer an attractive alternative by providing derivative-free optimization that is inherently well-suited to parallel computation. This makes them particularly attractive for trajectory optimization problems with non-smooth dynamics. However, naive sampling approaches, such as random search, suffer from poor sampling efficiency a slow convergence rate, and may be unable to converge to accurate solutions. Advances in sampling-based control over the past few years, particularly Model Predictive Path Integral (MPPI) control [theodorou2017mppi], despite their relative simplicity, have demonstrated promising results by incorporating more advanced sampling strategies and trajectory averaging schemes.
@@
In this paper, we present a reference-free robotic control framework that challenges the paradigm requiring hand-crafted locomotion patterns. While recently introduced sampling-based MPC methods rely on gait references, whether through Raibert heuristics [alvarezpadilla2024realtimewholebodycontrollegged], foot swing trajectories [Turrisi2024OnTB], or predefined contact sequences [xue2024fullordersamplingbasedmpctorquelevel], our approach enables the discovery of emergent locomotion purely from high-level goal specifications within a cost function. In summary, the key contributions are:
@@
  • A sampling-based MPC framework that enables reference-free motion discovery, without reliance on gait priors or offline pre-training.
@@
  • A dual-space spline parameterization that jointly samples position and velocity control points, improving exploration and dynamic consistency. Constraining endpoint derivatives further allows for bound-preserving rules that prevent overshooting joint limits.
@@ II Related Work
TABLE I: Comparison of sampling-based MPC for legged robot control. We combine the noise annealing strategy with the computational efficiency of prior work, while fully eliminating the need for reference gait requirements.
Several works have explored sampling-based methods for robotic control. A common strategy is to refine a nominal action sequence iteratively, typically parameterized as a spline, using random perturbations. For instance, Howell et al. [howell2022predictivesamplingrealtimebehaviour] evaluate predictive sampling (PS) as a zero-order baseline for comparison against methods like vanilla gradient descent, the iterative Linear Quadratic Regulator (iLQR) [Li2004IterativeLQ] and MPPI [theodorou2017mppi]. Despite its simplicity, PS achieves competitive performance and supports real-time tuning. The reported experiments, however, were restricted to torque-space splines in the MuJoCo [Todorov12mujoco] simulator and required workstation-class hardware.
More informative updates arise from leveraging statistics across all sampled trajectories, as done in MPPI. This approach has shown strong performance in real-time control for racing, where parallel rollouts can be combined efficiently [williams2016aggressive_driving]. Turrisi et al. [Turrisi2024OnTB] demonstrated MPPI on a 12-DOF quadruped using GPU-accelerated rigid-body dynamics, achieving 10k rollouts for a 12-step horizon in under $20\text{\,}\mathrm{ms}$, but with gait frequencies fixed a priori. Extensions to MPPI include DIAL-MPC [xue2024fullordersamplingbasedmpctorquelevel], which introduces a two-level annealing schedule inspired by diffusion models [pan2024modelbaseddiffusiontrajectoryoptimization]. This method broadens exploration in early iterations and later horizon steps, while gradually refining actions closer to execution. Evaluation relies on GPU-based simulation with thousands of parallel rollouts and reference gait tracking. Building on DIAL-MPC, Crestaz et al.  [crestaz2025td] add a constraint-enforcing mechanisms and a terminal value function approximation for longer-horizon reasoning.
Recent work has also investigated CPU-based implementations. Alvarez-Padilla et al. [alvarezpadilla2024realtimewholebodycontrollegged] propose real-time MPPI in joint space at $100\text{\,}\mathrm{Hz}$, with torques generated via a PD controller at $20\text{\,}\mathrm{kHz}$. Their system rolls out 30–50 trajectories in MuJoCo [Todorov12mujoco], avoiding fixed contact sequences but still relying on reference heuristics for robust behavior.
In contrast to these approaches, our method emphasizes low-sample efficiency and generality: it operates entirely on CPU with a limited number of rollouts, does not rely on gait priors, and employs a dual-space spline parameterization that samples both position and velocity targets for the PD controller. Unlike prior joint-space sampling methods, which provide only position references (and therefore set velocity targets to zero), our formulation yields dynamically consistent PD targets and richer exploration.
@@ III Methodology
We formulate our controller within the MPPI framework [theodorou2017mppi], which provides a sampling-based approach to trajectory optimization in a receding-horizon loop. At each control step, MPPI maintains a nominal control sequence over a finite horizon, perturbs this sequence with Gaussian noise, evaluates the resulting trajectories under a cost objective, and updates the nominal sequence via importance-weighted averaging before executing the first control input.
Building on this foundation and using a diffusion-inspired annealing scheme that structures noise, we introduce two practical enhancements: (i) a dual-space spline parameterization that jointly optimizes position and velocity control points to improve exploration and enforce dynamically consistent trajectories, (ii) reference-free cost formulations that support emergent locomotion without gait priors. To further improve stability and convergence, we integrate state prediction and warm-starting strategies to maintain temporal consistency across optimization steps. A complete overview of the framework is shown in Fig. 1(a) and Tab. I presents a comparison with recent MPPI-based methods.
@@ III-A Search-space parametrization
Defining the search space is crucial for effective random-search optimization. Unlike methods that sample in torque space, which typically require extensive parallel simulations on GPUs [vlahov2024mppi], we adopt joint-space sampling with a subsequent PD controller for impedance control, transforming joint-space references into torque references. By delegating low-level torque control to the PD controller, our method achieves stable configurations efficiently. This choice enables us to leverage efficient CPU-based simulation at high frame rates, requiring only 20-30 parallel trajectories to control a quadruped robot, and up to 60-70 parallel trajectories for a humanoid robot, compared to GPU-based simulators that offer high throughput but slower per-frame execution.
Dual-space spline parameterization. Control sequences are parameterized using cubic Hermite splines defined by $K$ node control points per controlled degree of freedom. Each spline node consist of both a position and a velocity component,
$\theta_{k}=(\theta_{k}^{q},\theta_{k}^{v}),$
so that the continuous joint trajectory $q(t)$ and its time derivative are reconstructed via Hermite interpolation. Denoting the set of spline basis functions by $\mathcal{H}_{k}^{(3)}(t)$ and their time derivatives by $\dot{\mathcal{H}}_{k}^{(3)}(t)$, the trajectory reads
$q(t)\;=\;\sum_{k=0}^{K-1}\theta_{k}^{q}\,\mathcal{H}_{k}^{(3)}(t)\;+\;\theta_{k}^{v}\,\dot{\mathcal{H}}_{k}^{(3)}(t).$ (1)
Figure 3: Plot comparison of different spline types with the same interpolation points (red), resulting in different normalized position and velocity trajectories. Cubic Hermite splines exhibit a lower variance than quadratic and cubic splines, resulting in finer sampling granularity.
By explicitly controlling position (joint position $q$) and its first-order time derivatives (joint velocity $v$) at the spline control points, Hermite splines provide more flexible representations of motion and enable better coverage of the trajectory space, as illustrated in Fig. 3. To prevent spline interpolants from crossing joint limits, we limit per-node derivatives based on the distance to the nearest bound. For node value $\theta_{k}^{q}\in[q_{min},q_{max}]$ and node spacing $\Delta t$, we impose
@@
This inexpensive clamp effectively suppresses mid-interval excursions. In Fig. 3, the red dashed lines indicate the upper bounds. While quadratic and cubic position splines frequently overshoot these limits, our cubic Hermite formulation respects the bounds throughout the trajectory. It is worth noting that quadratic or cubic splines inherently limit expressiveness: if the initial velocity is low, both do not generate significant mid-trajectory velocity changes. In contrast, Hermite splines naturally incorporate velocity information at control points, exhibit lower variance than quadratic and cubic splines, and result in finer sampling granularity. This richer parameterization enables a greater diversity of behaviors to be explored during sampling-based optimization, as shown in our experiments in Section IV.
@@ III-B Noise annealing
The annealing schedule is motivated by the iterative nature of receding-horizon MPC and follows the diffusion-inspired design introduced in DIAL-MPC [xue2024fullordersamplingbasedmpctorquelevel]. A nominal control sequence is refined over $I$ internal iterations at each time step before executing the first action, and then the horizon rolls forward. Consequently, control decisions farther in the horizon receive more updates (and thus can be explored more aggressively) while controls near execution should be more stable, with less variance.
Trajectory-level annealing. We therefore reduce exploration variance across iterations, transitioning from exploration to exploitation as the nominal trajectory is refined. For spline control points, this corresponds to shrinking the sampling covariance over iterations $i=I,\dots,1$:
@@
Action-level annealing. We additionally increase exploration for control points that correspond to later actions in the horizon. For spline node index $k\in\{0,\dots,K-1\}$, this yields
@@
III-C Algorithm description
The algorithm maintains a nominal sequence of spline control points in joint position $\theta^{q}$ and velocity $\theta^{v}$, defining a smooth trajectory over the prediction horizon via cubic Hermite interpolation. The position control points are initialized from a stable standing configuration, while velocity points start at zero. At each control step, new perturbed control points $(\theta_{n}^{q},\theta_{n}^{v})$ are sampled around the nominal points with structured noise from the annealing scheme. Interpolating these perturbed points yields a batch of candidate trajectories $\tau_{n}=(\tau_{n}^{q},\tau_{n}^{v})$, which are rolled out in Simple [carpentier2024compliantrigidcontactsimulation], evaluated under the task cost, and assigned importance weights. The nominal sequence is then updated by importance-weighted averaging, and new control points are extracted by resampling the updated trajectory at spline node times, distributed uniformly across the horizon.
The exponential weighting scheme considers the relative quality of all candidates, see Eq. (6), and enables effective trajectory synthesis. The softmax smoothing naturally concentrates probability mass around high-quality solutions and can be interpreted as performing approximate natural gradient descent on a smoothed surrogate function [jordana2025introductionzeroorderoptimizationtechniques]. The normalized weight $\omega_{n}$ for each trajectory $n$ is computed as
where $S_{n}$ is the cost of trajectory $n$ and $\lambda$ is a temperature parameter that controls the selectiveness of the weighting. In practice, we implement the calculation of $\omega_{n}$ using min-max-normalization of costs [JMLR:v11:theodorou10a]:
and the weights are computed as
Overall, we build on MPPI with noise annealing and extend it with two key enhancements that enable reference-free locomotion discovery with a low number of samples: (1) dual-space spline sampling, which jointly perturbs position and velocity control points $(\theta^{q},\theta^{v})$ with physics-aware scaling, and (2) best trajectory tracking, which separates the evolving nominal sequence $\tau_{0}$ from the executed actions $\tau_{best}$ for robustness and as a safeguard for consistent performance across iterations. Alg. 1 summarizes our complete approach, integrating these aspects within the MPPI framework.

### v2

@@ Abstract
We present a sampling-based model predictive control (MPC) framework that enables emergent locomotion without relying on handcrafted gait patterns or predefined contact sequences. Our method discovers diverse motion patterns, ranging from trotting to galloping, robust standing policies, jumping, and handstand balancing, purely through the optimization of high-level objectives. Building on model predictive path integral (MPPI), we propose a cubic Hermite spline parameterization that operates on position and velocity control points. Our approach enables contact-making and contact-breaking strategies that adapt automatically to task requirements, requiring only a limited number of sampled trajectories. This sample efficiency enables real-time control on standard CPU hardware, eliminating the GPU acceleration typically required by other state-of-the-art MPPI methods. We validate our approach on the Go2 quadrupedal robot, demonstrating a range of emergent gaits and basic jumping capabilities. In simulation, we further showcase more complex behaviors, such as backflips, dynamic handstand balancing and locomotion on a Humanoid, all without requiring reference tracking or offline pre-training.
@@ I Introduction
(a) Overview of the framework showing the Hermite spline parametrization, noise schedule and reference-free costs.
@@
Figure 1: Our reference-free sampling-based MPC framework enables emergent jumping motion experimentally achieved on the Go2 robot without any guiding reference.
@@
Online robot control presents a trade-off between solution quality and computational efficiency. While learning-based methods such as reinforcement learning (RL) can generate impressive movements on complex robots, they require extensive offline training. They may also fail to generalize or adapt to new environments or tasks not seen during training [16]. Despite ongoing efforts to improve sample efficiency [17], RL remains orders of magnitude too slow for online adaptation, and simplified models used to accelerate training can hinder sim-to-real transfer [8, 29]. Typically, RL methods rely heavily on engineered rewards, such as phase clocks, air-time penalties, or foot-clearance objectives, to enforce structured contact behavior [12, 2, 9].
@@
On the other hand, trajectory optimization (TO) relies on derivative information to find high-quality solutions, but it requires accurate and informative gradients. This may be unavailable or unreliable in contact-rich scenarios, leading most TO methods to use predefined contact sequences [4, 13]. Contact-implicit TO removes the need for predefined contact schedules [21] and has enabled advanced demonstrations of complex whole-body behaviors [19, 15]. However, many formulations rely on simplified contact models or approximations to remain tractable, which can introduce a mismatch with real dynamics and produce only approximate solutions in practice. These methods also necessitate hand-crafted cost functions to obtain good contact sequences [15].
@@
Sampling-based methods offer an attractive alternative by providing derivative-free optimization that is well-suited to parallel computation. This makes them particularly attractive for trajectory optimization problems with non-smooth dynamics. However, naive sampling approaches, such as random search, suffer from poor sampling efficiency and slow convergence, and may fail to converge to accurate solutions. Advances in sampling-based control over the past few years, particularly Model Predictive Path Integral (MPPI) control [26], despite their relative simplicity, have demonstrated promising results by incorporating more advanced sampling strategies and trajectory averaging schemes.
@@
In this paper, we present a reference-free robotic control framework that challenges the paradigm requiring hand-crafted locomotion patterns. While recently introduced sampling-based MPC methods rely on gait references, whether through Raibert heuristics [1], foot swing trajectories [24], or predefined contact sequences [28], our approach enables the discovery of emergent locomotion purely from high-level goal specifications within a cost function. In summary, the key contributions are:
@@
  • A sampling-based MPC framework that enables reference-free motion discovery without reliance on gait priors or offline pre-training.
@@
  • A cubic Hermite spline parameterization that jointly samples position and velocity control points to improve exploration and dynamic consistency. Constraining endpoint derivatives allows for bound-preserving rules that prevent overshooting joint limits.
@@ II Related Work
TABLE I: Comparison of sampling-based MPC for legged robot control. We combine the noise-annealing strategy with the computational efficiency of prior work while eliminating the need for reference gait requirements.
Several works have explored sampling-based methods for robotic control. A common strategy is to refine a nominal action sequence iteratively, typically parameterized as a spline, using random perturbations. Howell et al. [11] evaluate predictive sampling (PS) as a zero-order baseline for comparison against methods like gradient descent, the iterative Linear Quadratic Regulator (iLQR) [18] and MPPI [26]. Despite its simplicity, PS achieves competitive performance and supports real-time tuning. The reported experiments, however, were restricted to torque-space splines in the MuJoCo [23] simulator and required workstation-class hardware.
More informative updates arise from leveraging statistics across all sampled trajectories, as in MPPI. This has shown strong performance in real-time control for racing, where parallel rollouts can be combined efficiently [27]. Related work has explored improving MPPI through structured proposal distributions, for example, via feature-based sampling [10]. Turrisi et al. [24] demonstrated MPPI on a 12-DOF quadruped using GPU-accelerated rigid-body dynamics, achieving 10k rollouts for a 12-step horizon in under $20\text{\,}\mathrm{ms}$, but with gait frequencies fixed a priori. Extensions to MPPI include DIAL-MPC [28], which introduces a two-level annealing schedule inspired by diffusion models [20]. This method broadens exploration in early iterations and later horizon steps, while gradually refining actions closer to execution. Evaluation relies on GPU-based simulation with thousands of parallel rollouts and reference gait tracking. Building on DIAL-MPC, Crestaz et al.  [6] add a constraint-enforcing mechanism and a terminal value function approximation for longer-horizon reasoning.
Recent work has also investigated CPU-based implementations. Alvarez-Padilla et al. [1] propose real-time MPPI in joint space at $100\text{\,}\mathrm{Hz}$, with torques generated via a PD controller at $20\text{\,}\mathrm{kHz}$. Their system rolls out 30–50 trajectories in MuJoCo [23], avoiding fixed contact sequences while still relying on reference heuristics for robust behavior.
In contrast to these approaches, our method emphasizes low-sample efficiency and generality. It operates entirely on the CPU, uses a limited number of rollouts, does not rely on gait priors, and employs cubic Hermite splines to sample both position and velocity targets for the PD controller. Unlike prior joint-space sampling methods, which provide only position references (with velocity targets set to zero), our formulation yields dynamically consistent PD targets and enables richer exploration.
@@ III Methodology
We formulate our controller within the MPPI framework [26], which provides a sampling-based approach to trajectory optimization in a receding-horizon loop. At each control step, MPPI maintains a nominal control sequence over a finite horizon, perturbs this sequence with Gaussian noise, evaluates the resulting trajectories under a cost objective, and updates the nominal sequence via importance-weighted averaging before executing the first control input.
Building on this foundation and using a diffusion-inspired annealing scheme that structures noise, we introduce several practical enhancements. We use a cubic Hermite spline parameterization that jointly optimizes position and velocity control points to improve exploration and enforce dynamically consistent trajectories, and a reference-free cost formulation that supports emergent locomotion without gait priors. To further improve stability and convergence, we integrate state prediction and warm-starting strategies to maintain temporal consistency across optimization steps. A complete overview of the framework is shown in Fig. 1(a) and Tab. I presents a comparison with recent MPPI-based methods.
@@ III-A Search-space parametrization
Defining the search space is critical for effective random-search optimization. In contrast to approaches that sample directly in torque space, often together with massively parallel GPU simulation [25], we sample reference trajectories in joint space and track them using a PD controller that maps joint-space references to torque commands. This decouples high-level trajectory exploration from low-level torque regulation, enabling efficient CPU-based simulation at high frame rates. In our implementation, this requires only 20–30 parallel trajectories for quadruped control and 60–70 for humanoid control.
Cubic Hermite spline parameterization. For each actuated degree of freedom, the control sequence is parameterized by a cubic Hermite spline with $K$ nodes. The optimization variables are the node parameters
$\theta_{k}=(\theta_{k}^{q},\theta_{k}^{v}),\qquad k=0,\dots,K-1,$
where $\theta_{k}^{q}=q(t_{k})$ and $\theta_{k}^{v}=\dot{q}(t_{k})$ denote the joint position and joint velocity at node $k$. Throughout the paper, the superscript v denotes joint-velocity quantities associated with $q$. The nodes are uniformly spaced in time with interval $\Delta t$, and we consider now one spline segment over $t\in[t_{k},t_{k+1}]$. With the normalized local time
$s=\frac{t-t_{k}}{\Delta t}\in[0,1],$ (1)
the joint trajectory is reconstructed using the standard cubic Hermite interpolant [7]
$q(t)$ $=h_{00}(s)\,\theta_{k}^{q}\;\;\>+h_{10}(s)\,\Delta t\,\theta_{k}^{v}$ (2)
$+h_{01}(s)\,\theta_{k+1}^{q}+h_{11}(s)\,\Delta t\,\theta_{k+1}^{v},$
$\text{where}\quad h_{00}(s)$ $=2s^{3}-3s^{2}+1,$ $h_{10}(s)$ $=s^{3}-2s^{2}+s,$ (3)
$h_{01}(s)$ $=-2s^{3}+3s^{2},$ $h_{11}(s)$ $=s^{3}-s^{2}$ (4)
are the cubic Hermite basis functions. This parameterization is applied only to the actuated joints. The floating-base motion is not parameterized directly. Instead, it emerges from the forward dynamics rollout under the sampled joint references and contact interactions.
Figure 3: Comparison of spline parameterizations for the same interpolation points (red). Quadratic, cubic, and cubic Hermite splines produce different position and velocity profiles. By parameterizing both node positions and velocities, cubic Hermite splines provide additional local shape control. Derivative clamping reduces overshoot near joint limits.
Compared with position-only spline parameterizations, cubic Hermite splines control both node positions and node velocities. This provides additional local control over the shape of the generated reference trajectories, as illustrated in Fig. 3. To reduce overshoot near joint limits, we limit per-node derivatives based on the distance to the nearest bound. For node value $\theta_{k}^{q}\in[q_{min},q_{max}]$ and node spacing $\Delta t$, we impose
@@
This inexpensive clamp reduces mid-interval excursions and improves preservation of bounds in practice. In Fig. 3, the red dashed lines indicate the upper position bounds. While quadratic and cubic position splines often overshoot these bounds, our cubic Hermite formulation respects them throughout the trajectory. In the current formulation, the spline parameterization enforces position bounds via the derivative clamp above, but does not explicitly enforce joint-velocity limits at any intermediate time. In practice, feasibility is further shaped by the rollout dynamics, PD tracking, and the cost function.
@@ III-B Noise annealing
The annealing schedule is motivated by the iterative nature of receding-horizon MPC and follows the diffusion-inspired design introduced in DIAL-MPC [28]. A nominal control sequence is refined over $I$ internal iterations at each time step before executing the first action, after which the horizon advances. Consequently, control decisions farther in the horizon receive more updates (and thus can be explored more aggressively) while controls near execution should be more stable, with less variance.
Trajectory-level annealing. We therefore reduce the variance of exploration across iterations, transitioning from exploration to exploitation as the nominal trajectory is refined. For spline control points, this corresponds to shrinking the sampling covariance over iterations $i=I,\dots,1$:
@@
Action-level annealing. We also increase exploration of control points corresponding to later actions in the horizon. For spline node index $k\in\{0,\dots,K-1\}$, this yields
@@
III-C Algorithm description
The algorithm maintains a nominal sequence of spline control points in joint position $\theta^{q}$ and velocity $\theta^{v}$, defining a smooth trajectory over the prediction horizon via cubic Hermite interpolation. The position control points are initialized from a stable standing configuration and the velocity control points are initialized to zero. At each control step, new perturbed control points $(\theta_{n}^{q},\theta_{n}^{v})$ are sampled around the nominal points with structured noise from the annealing scheme. Interpolating these perturbed points yields a batch of candidate trajectories $\tau_{n}=(\tau_{n}^{q},\tau_{n}^{v})$, which are rolled out in Simple [5], evaluated under the task cost, and assigned importance weights. The nominal sequence is then updated via importance-weighted averaging, and new control points are extracted by resampling the updated trajectory at spline node times, uniformly distributed across the horizon.
The exponential weighting scheme considers the relative quality of all candidates (see Eq. (9)) and enables effective trajectory synthesis. The softmax smoothing concentrates probability mass around high-quality solutions and can be interpreted as performing approximate natural gradient descent on a smoothed surrogate function [14]. The normalized weight $\omega_{n}$ for each trajectory $n$ is computed as
where $S_{n}$ is the cost of trajectory $n$ and $\lambda$ is a temperature parameter that controls the selectiveness of the weighting. In practice, we implement the calculation of $\omega_{n}$ using min-max-normalization of costs [22]:
and the weights are computed as
Overall, we build on MPPI with noise annealing and extend it with two key enhancements that enable reference-free locomotion discovery with a low number of samples: (1) cubic Hermite spline sampling, which jointly perturbs position and velocity control points $(\theta^{q},\theta^{v})$ with physics-aware scaling, and (2) best trajectory tracking, which separates the evolving nominal sequence $\tau_{0}$ from the executed actions $\tau_{best}$ for robustness and as a safeguard for consistent performance across iterations. Alg. 1 summarizes our complete approach, integrating these aspects within the MPPI framework.